# GENERator: A Long-Context Generative Genomic Foundation Model

### v4

@@ Abstract
The rapid advancement of DNA sequencing has produced vast genomic datasets, yet the interpretation and rational engineering of sequence function remain fundamental challenges. Recent large language models (LLMs) have opened new avenues for genomic analysis; however, existing approaches are frequently constrained by limited training scope, restricted generative flexibility, or prohibitive computational cost. In this study, we introduce GENERator, a generative genomic foundation model designed for long-context DNA modeling, with a context length of 98k nucleotides, pre-trained on 386 billion nucleotides of eukaryotic DNA. GENERator demonstrates strong intrinsic capabilities arising directly from pre-training. Unsupervised embedding analyses reveal latent organization consistent with phylogenetic relationships. Sequence recovery benchmarks show that GENERator achieves generative accuracy matching or exceeding state-of-the-art baselines with substantially improved computational efficiency. In a zero-shot setting, GENERator further provides competitive variant effect prediction performance relative to alignment-based methods, while remaining fully alignment-free and broadly applicable across species. Beyond training-free evaluation, GENERator consistently delivers strong performance through task-specific fine-tuning on established genomic benchmarks. We further demonstrate practical generative applications enabled by the model. GENERator can generate protein-coding DNA sequences that translate into structurally plausible proteins and, through a prompt-guided design framework, can generate cis-regulatory elements with targeted activity profiles, including synthetic super-enhancers validated by high-throughput UMI-STARR-seq assays. Collectively, these results establish GENERator as an efficient and biologically grounded foundation for genomic interpretation and programmable sequence design across diverse genomic contexts. Implementation details and supplementary resources are available at https://github.com/GenerTeam/GENERator.
@@ 1 Introduction
Beyond training-free evaluation, GENERator demonstrates strong adaptability through task-specific fine-tuning, achieving leading performance across established genomic benchmarks, including NT tasks [21], Genomic Benchmarks [29], and the newly proposed Gener tasks. We further demonstrate its utility in biologically grounded generative applications across multiple biological scales. Consistent with the central dogma of molecular biology [19], GENERator generates protein-coding DNA sequences that translate into structurally plausible proteins. Moreover, through a prompt-guided design framework, the model enables programmable cis-regulatory element (CRE) design, producing regulatory sequences with targeted activity profiles, including synthetic super-enhancers that exceed the activity of natural counterparts.
@@ 2.1 Embedding Clustering
Figure 2: Comprehensive benchmarking of GENERator. (A) Sequence recovery accuracy is compared across representative DNA foundation models. GENERator and Evo2 substantially outperform all other baseline models. At matched parameter scales, GENERator-1B consistently exceeds Evo2-1B in overall performance while achieving markedly higher generation efficiency. Scaling up to GENERator-3B yields clear performance gains over the 1B model and approaches the performance of Evo2-7B. (B) Generation time is measured on a single L40S GPU by conditioning on a 6k input sequence and generating the next 50 bp. GENERator exhibits tens-fold faster generation speed compared to Evo2 across model sizes, highlighting its substantially improved computational efficiency. (C) Sequence recovery performance is evaluated under identical model architectures, batch sizes, and training steps across six taxonomic groups. The 6-mer tokenizer achieves the best overall performance, whereas BPE tokenization performs consistently worse, likely due to the hierarchical nature of the BPE vocabulary. (D) Performance on ClinVar variant effect prediction is evaluated using AUPRC and AUROC. GENERator and Evo2 significantly outperform other sequence-based self-supervised models. GENERator-1B consistently surpasses Evo2-1B, while GENERator-3B achieves performance comparable to Evo2-7B. Although MSA-based methods such as GPN-MSA and CADD retain advantages on this task, GENERator and Evo2 operate without MSA information, enabling straightforward application to non-model organisms. (E) Comparison across diverse supervised benchmarks, including Revised NT tasks, Original NT tasks, Genomic Benchmarks, and Gene tasks, shows that GENERator achieves the strongest overall performance across all evaluated settings. (F-G) Central-dogma tasks are evaluated on the cytochrome P450 family (F) and histone family (G). Generated DNA sequences exhibit stable and continuous coding structures. Translated proteins show low perplexity under the protein language model ProGen2, indicating high protein-likeness. Structural validation using AlphaFold3 followed by Foldseek reveals that most generated proteins fall within high-confidence regions. Importantly, many generated proteins display strong structural similarity to natural counterparts despite low sequence identity (<0.3<0.3), demonstrating that GENERator captures protein grammar rather than memorizing natural sequences.
@@ 2.4 Benchmark Evaluations
Given the recent revision of the NT task dataset [21], we evaluated models on both the original and revised benchmark sets. The average performance across tasks in each benchmark is shown in the left and middle panels of Figure 2E, with per-task results reported in Tables S3 and S4. Across both versions, GENERator exhibits the strongest average performance among evaluated models. Notably, the reduced performance observed for GENERator-All highlights the importance of the functional sequence training strategy (Section 4.1). Enformer maintains competitive performance on chromatin profile and regulatory element tasks, reflecting its original supervised training focus on these domains [4]. The recently released Nucleotide Transformer v2 (NT-v2), despite its smaller size (500M parameters), achieves higher average performance than the original NT model (2.5B parameters). In contrast, models employing BPE tokenization (DNABERT-2, GROVER) or single-nucleotide tokenization (HyenaDNA, Caduceus) do not show consistent advantages on these benchmarks.
@@
We further evaluated models on Genomic Benchmarks [29], which primarily target human genomic tasks, and on the newly proposed Gener tasks, designed to assess gene type and taxonomic classification across diverse organisms and sequence lengths. Average performance across all tasks in both benchmark suites is summarized in the right panel of Figure 2E, with detailed results reported in Tables S5 and S6.
@@ 2.5 Central Dogma
As shown in Figure 2F for cytochrome P450 and Figure 2G for histone, the length distributions of both generated DNA sequences and their translated protein products closely match those of natural sequences after deduplication. Critically, 99.9% of generated sequences were successfully translated into proteins using standard codon tables, indicating stable coding structure without premature stop codons or frameshift errors. This high translation success rate demonstrates that the model preserves essential constraints of protein-coding regions.
@@ 2.6 Cis-Regulatory Element Design with High-throughput Validation
Collectively, these findings demonstrate that GENERator is capable of exploring regulatory sequence space beyond naturally occurring cis-regulatory elements. The identification of both super-enhancers and ultra-low-activity silencers highlights the potential of generative genomic language models to not only reproduce, but also expand the functional repertoire of regulatory genomics.
@@ 4.5.3 Cis-Regulatory Element Design
All Python scripts used for downstream analyses and figure generation are available on GitHub at https://github.com/GenerTeam/GENERator.
@@ Appendix G Supplementary Figures & Tables
Figure S4: Comparative sequence recovery accuracy across tokenization strategies and model architectures. Each cell shows accuracy (y-axis) versus input tokens (x-axis, 8-1024) for a taxonomic group and prediction length (1-30 bp). Lines represent: BPE tokenizers (all vocabulary sizes, consistently poor), K-mer tokenizers with K=1-2 (moderate performance), K-mer tokenizers with K=3-8 (superior performance), and Mamba architectures (Mamba×1: same tokens as 1-mer; Mamba×6: same bp as 6-mer). While K=3-8 perform similarly well, 6-mer achieves optimal balance - smaller K values (K=3-5) provide limited context extension, while larger K (7,8) suffer in data-scarce groups (fungi, protozoa) and with fewer tokens. Mamba architectures reveal critical insights: both Mamba×1 and Mamba×6 achieve only K=1-level accuracy despite the 6-fold longer context in Mamba×6. Mamba×6 shows initial advantage with few tokens (effectively more tokens) but is surpassed as token count increases. The convergence of Mamba×1 (1024 tokens) and Mamba×6 (6144 tokens) at similar accuracy indicates that while SSMs can process long inputs, they lack proportional gains in contextual understanding compared to transformers.
Figure S5: Comparative sequence recovery accuracy across genomic foundation models. Each cell displays accuracy (y-axis) versus input tokens (x-axis, 8-1024) for a taxonomic group and prediction length (1-30 bp). Lines represent various baseline models: MLM models (consistently poor as expected due to lack of generative capability); HyenaDNA (autoregressive but severely limited by its 55M parameters, a capacity highly disproportionate to its 1M context length); and the leading autoregressive models GENERator and Evo2. We compare two Evo2 configurations: [Evo2]×1 (same token count as other models, representing comparable efficiency) and [Evo2]×6 (same nucleotide count as GENERator, but with inference times several to dozens of times slower than GENERator). Evo2 architectures, also based on the SSM framework, reveal patterns similar to Mamba-2 in Figure S4: [Evo2]×6 shows initial advantage with few tokens (more effective tokens) but this advantage diminishes rapidly within a few hundred tokens. Critically, [Evo2]×1 (1024 tokens) and [Evo2]×6 (6144 tokens) achieve nearly identical accuracy, indicating ineffective utilization of extended context despite the long-context capabilities claimed by Evo2. While Evo2-7B achieves higher absolute accuracy (attributable to larger training data, more steps, and greater parameters), GENERator demonstrates superior parameter efficiency: GENERator-1B surpasses Evo2-1B, and GENERator-3B approaches Evo2-7B performance. Notably, GENERator maintains a clear upward accuracy trend even at the maximum tested length (1024 tokens, 6144 bp), suggesting better long-context utilization.
@@
Figure S7: Comparison of pre-training loss across tokenization strategies. Loss curves are shown for various tokenizers: BPE (vocabulary sizes 512-8192), K-mer (K=1-8), and Mamba+1mer. The substantial variation in loss values—ranging from roughly 1 for 1-mer (4-way classification) to significantly higher for 8-mer (65536-way classification)—demonstrates the inherent incomparability of pre-training loss across different tokenization schemes. This fundamental limitation motivated the development of our sequence recovery task (Section 2.2), which enables fair cross-tokenizer and cross-architecture comparison. Close examination reveals why BPE underperforms: in the BPE-512 vs. 4-mer comparison, 4-mer initially shows higher loss but narrows the gap with training; similarly, in BPE-4096 vs. 5-mer, 5-mer exhibits faster loss reduction over time. These patterns indicate that BPE tokenizers present greater learning challenges, supporting our hypothesis that next token prediction pre-training may not be optimally aligned with BPE tokenization for genomic sequences. This misalignment stems from the hierarchical nature of BPE vocabulary: unlike human language with clear lexical boundaries, DNA sequences lack explicit delimiters, creating nested relationships among BPE tokens. For example, when predicting target token GCCT, partially correct predictions like G, GC, or GCC are penalized as incorrect, unnecessarily complicating the training process. The Mamba+1mer model exhibits initial training instability, likely due to the extreme sequence length of 98k tokens, but gradually stabilizing and converging to a loss level comparable to 1-mer, consistent with its sequence recovery performance.

### v5

@@ Abstract
The rapid advancement of DNA sequencing has produced vast genomic datasets, yet the interpretation and rational engineering of sequence function remain fundamental challenges. Recent large language models (LLMs) have opened new avenues for genomic analysis; however, existing approaches are frequently constrained by limited training scope, restricted generative flexibility, or prohibitive computational cost. In this study, we introduce GENERator, a generative genomic foundation model designed for long-context DNA modeling, with a context length of 98k nucleotides, pre-trained on 386 billion nucleotides of eukaryotic DNA. GENERator demonstrates strong intrinsic capabilities arising directly from pre-training. Unsupervised embedding analyses reveal latent organization consistent with phylogenetic relationships. Sequence recovery benchmarks show that GENERator achieves generative accuracy matching or exceeding state-of-the-art baselines with substantially improved computational efficiency. In a zero-shot setting, GENERator further provides competitive variant effect prediction performance relative to alignment-based methods, while remaining fully alignment-free and broadly applicable across species. Beyond training-free evaluation, GENERator consistently delivers strong performance through task-specific fine-tuning on established genomic benchmarks. We further demonstrate practical generative applications enabled by the model. GENERator can generate protein-coding DNA sequences that translate into structurally plausible proteins and, through a prompt-guided design framework, design cis-regulatory elements with targeted activity profiles, including synthetic enhancers whose regulatory strength exceeds that of natural genomic sequences, as validated by high-throughput UMI-STARR-seq assays. Collectively, these results establish GENERator as an efficient and biologically grounded foundation for genomic interpretation and programmable sequence design across diverse genomic contexts. Implementation details and supplementary resources are available at https://github.com/GenerTeam/GENERator.
@@ 1 Introduction
Beyond training-free evaluation, GENERator demonstrates strong adaptability through task-specific fine-tuning, achieving leading performance across established genomic benchmarks, including NT tasks [21], Genomic Benchmarks [29], and the newly proposed Gener tasks. We further demonstrate its utility in biologically grounded generative applications across multiple biological scales. Consistent with the central dogma of molecular biology [19], GENERator generates protein-coding DNA sequences that translate into structurally plausible proteins. Moreover, through a prompt-guided design framework, the model enables programmable cis-regulatory element (CRE) design, generating regulatory sequences with targeted activity profiles, including synthetic enhancers whose regulatory strength exceeds that observed in natural counterparts.
@@ 2.1 Embedding Clustering
Figure 2: Training-free evaluations of GENERator. (A) Sequence recovery accuracy is compared across representative DNA foundation models. GENERator and Evo2 substantially outperform all other baseline models. At matched parameter scales, GENERator-1B consistently exceeds Evo2-1B in overall performance while achieving markedly higher generation efficiency. Scaling up to GENERator-3B yields clear performance gains over the 1B model and approaches the performance of Evo2-7B. (B) Generation time is measured on a single L40S GPU by conditioning on a 6k input sequence and generating the next 50 bp. GENERator exhibits tens-fold faster generation speed compared to Evo2 across model sizes, highlighting its substantially improved computational efficiency. (C) Sequence recovery performance is evaluated under identical model architectures, batch sizes, and training steps across six taxonomic groups. The 6-mer tokenizer achieves the best overall performance, whereas BPE tokenization performs consistently worse, likely due to the hierarchical nature of the BPE vocabulary. (D) Performance on ClinVar variant effect prediction is evaluated using AUPRC and AUROC. GENERator and Evo2 significantly outperform other sequence-based self-supervised models. GENERator-1B consistently surpasses Evo2-1B, while GENERator-3B achieves performance comparable to Evo2-7B. Although MSA-based methods such as GPN-MSA and CADD retain advantages on this task, GENERator and Evo2 operate without MSA information, enabling straightforward application to non-model organisms.
@@ 2.3 Variant Effect Prediction
Figure 3: Task-specific fine-tuning of GENERator. (A-C) Comparison across diverse supervised benchmarks, including Revised NT tasks (A), Original NT tasks (B), Genomic Benchmarks and Gene tasks (C), shows that GENERator achieves the strongest overall performance across all evaluated settings. (D-E) Central-dogma tasks are evaluated on the cytochrome P450 family (D) and histone family (E). Generated DNA sequences exhibit stable and continuous coding structures. Translated proteins show low perplexity under the protein language model ProGen2, indicating high protein-likeness. Structural validation using AlphaFold3 followed by Foldseek reveals that most generated proteins fall within high-confidence regions. Importantly, many generated proteins display strong structural similarity to natural counterparts despite low sequence identity (<0.3<0.3), demonstrating that GENERator captures protein grammar rather than memorizing natural sequences.
@@ 2.4 Benchmark Evaluations
Given the recent revision of the NT task dataset [21], we evaluated models on both the original and revised benchmark sets. The average performance across tasks in each benchmark is shown in Figure 3A and B, with per-task results reported in Tables S3 and S4. Across both versions, GENERator exhibits the strongest average performance among evaluated models. Notably, the reduced performance observed for GENERator-All highlights the importance of the functional sequence training strategy (Section 4.1). Enformer maintains competitive performance on chromatin profile and regulatory element tasks, reflecting its original supervised training focus on these domains [4]. The recently released Nucleotide Transformer v2 (NT-v2), despite its smaller size (500M parameters), achieves higher average performance than the original NT model (2.5B parameters). In contrast, models employing BPE tokenization (DNABERT-2, GROVER) or single-nucleotide tokenization (HyenaDNA, Caduceus) do not show consistent advantages on these benchmarks.
@@
We further evaluated models on Genomic Benchmarks [29], which primarily target human genomic tasks, and on the newly proposed Gener tasks, designed to assess gene type and taxonomic classification across diverse organisms and sequence lengths. Average performance across all tasks in both benchmark suites is summarized in the right panel of Figure 3C, with detailed results reported in Tables S5 and S6.
@@ 2.5 Central Dogma
As shown in Figure 3D for cytochrome P450 and Figure 3E for histone, the length distributions of both generated DNA sequences and their translated protein products closely match those of natural sequences after deduplication. Critically, 99.9% of generated sequences were successfully translated into proteins using standard codon tables, indicating stable coding structure without premature stop codons or frameshift errors. This high translation success rate demonstrates that the model preserves essential constraints of protein-coding regions.
@@ 2.6 Cis-Regulatory Element Design with High-throughput Validation
Collectively, these findings demonstrate that GENERator can explore regulatory sequence space beyond naturally occurring cis-regulatory elements. The identification of both high-activity enhancers and low-activity silencers whose regulatory strengths extend beyond the natural genomic range highlights the potential of generative genomic language models to not only reproduce, but also expand the functional repertoire of regulatory genomics.
@@ 4.5.3 Cis-Regulatory Element Design
Acknowledgements
The authors thank Guojie Zhang, Guangji Chen, Guangyong Chen, and Qinqin Long for their early support and helpful discussions during the initial stages of the project. We also thank Chuan Cao, Haiguang Liu, and Tao Qin for their support and insightful discussions during the final phase of the work. In addition, we thank Lulu Zhang for her assistance with administrative coordination and management throughout the project.
Author Contributions
Q.L. was the primary lead of the project and contributed to study conception, data preparation, model design, downstream analyses, and manuscript drafting and revision. W.W. made major technical contributions to model development, downstream evaluation, and open-source code implementation. Y.Z. performed CRE synthesis and validation experiments, conducted data analysis and visualization, and contributed to drafting the corresponding sections. Z.Z. carried out computational evaluations, particularly related to Evo2 benchmarking, and contributed to discussions.
Y.L. conceived and supervised the CRE design framework and oversaw the wet-lab experiments. R.C., J.Q., Y.B., and C.W. contributed to CRE-related data processing and analysis. M.L., K.F., Y.Zhu, and Z.Zhang contributed to discussions and manuscript drafting.
J.Y., F.F., and J.T. provided supervision, computational resources, and insightful discussions. H.X. and Z.W. provided overall project supervision, strategic guidance, and computational and personnel support. All authors reviewed and approved the final manuscript.
Competing Interests
The authors declare no competing interests.
@@ Acknowledgements
Code Availability
All Python scripts used for downstream analyses are available on GitHub at https://github.com/GenerTeam.
@@ Appendix G Supplementary Figures & Tables
Figure S4: Comparative sequence recovery accuracy across tokenization strategies and model architectures. Each cell shows accuracy (y-axis) versus input tokens (x-axis, 8-1024) for a taxonomic group and prediction length (1-30 bp). Lines represent: BPE tokenizers (all vocabulary sizes, consistently poor), $k$-mer tokenizers with kk=1-2 (moderate performance), $k$-mer tokenizers with kk=3-8 (superior performance), and Mamba architectures (Mamba×1: same tokens as 1-mer; Mamba×6: same bp as 6-mer). While kk=3-8 perform similarly well, 6-mer achieves optimal balance - smaller $k$ values (kk=3-5) provide limited context extension, while larger $k$ (7,8) suffer in data-scarce groups (fungi, protozoa) and with fewer tokens. Mamba architectures reveal critical insights: both Mamba×1 and Mamba×6 achieve only 1mer-level accuracy despite the 6-fold longer context in Mamba×6. Mamba×6 shows initial advantage with few tokens (effectively more tokens) but is surpassed as token count increases. The convergence of Mamba×1 (1024 tokens) and Mamba×6 (6144 tokens) at similar accuracy indicates that while SSMs can process long inputs, they lack proportional gains in contextual understanding compared to transformers.
Figure S5: Comparative sequence recovery accuracy across genomic foundation models. Each cell displays accuracy (y-axis) versus input tokens (x-axis, 8-1024) for a taxonomic group and prediction length (1-30 bp). Lines represent various baseline models: MLM models (consistently poor as expected due to lack of generative capability); HyenaDNA (autoregressive but severely limited by its 55M parameters, a capacity highly disproportionate to its 1M context length); and the leading autoregressive models GENERator and Evo2. We compare two Evo2 configurations: [Evo2]×1 (same token count as other models, representing comparable efficiency) and [Evo2]×6 (same nucleotide count as GENERator, but with inference times several to dozens of times slower than GENERator). Evo2 series reveal patterns similar to Mamba-2 in Figure S4: [Evo2]×6 shows initial advantage with few tokens (more effective tokens) but this advantage diminishes rapidly within a few hundred tokens. Critically, [Evo2]×1 (1024 tokens) and [Evo2]×6 (6144 tokens) achieve nearly identical accuracy, indicating ineffective utilization of extended context despite the long-context capabilities claimed by Evo2. While Evo2-7B achieves higher absolute accuracy (attributable to larger training data, more steps, and greater parameters), GENERator demonstrates superior parameter efficiency: GENERator-1B surpasses Evo2-1B, and GENERator-3B approaches Evo2-7B performance. Notably, GENERator maintains a clear upward accuracy trend even at the maximum tested length (1024 tokens, 6144 bp), suggesting better long-context utilization.
@@
Figure S7: Comparison of pre-training loss across tokenization strategies. Loss curves are shown for various tokenizers: BPE (vocabulary sizes 512-8192), $k$-mer (kk=1-8), and Mamba+1mer. The substantial variation in loss values—ranging from roughly 1 for 1-mer (4-way classification) to significantly higher for 8-mer (65536-way classification)—demonstrates the inherent incomparability of pre-training loss across different tokenization schemes. This fundamental limitation motivated the development of our sequence recovery task (Section 2.2), which enables fair cross-tokenizer and cross-architecture comparison. Close examination reveals why BPE underperforms: in the BPE-512 vs. 4-mer comparison, 4-mer initially shows higher loss but narrows the gap with training; similarly, in BPE-4096 vs. 5-mer, 5-mer exhibits faster loss reduction over time. These patterns indicate that BPE tokenizers present greater learning challenges, supporting our hypothesis that next token prediction pre-training may not be optimally aligned with BPE tokenization for genomic sequences. This misalignment stems from the hierarchical nature of BPE vocabulary: unlike human language with clear lexical boundaries, DNA sequences lack explicit delimiters, creating nested relationships among BPE tokens. For example, when predicting target token GCCT, partially correct predictions like G, GC, or GCC are penalized as incorrect, unnecessarily complicating the training process. The Mamba+1mer model exhibits initial training instability, likely due to the extreme sequence length of 98k tokens, but gradually stabilizing and converging to a loss level comparable to 1-mer, consistent with its sequence recovery performance.